If an athlete is able to sprint 100 metres in 10 seconds, the average speed of the athlete is:


 


a. 36 km/h


b. 36 m/s


c. 6 km/h


d. 60 km/h


e. 38 km/h

Grade:11

3 Answers

Soumyadip
10 years ago

answer is A. 36 km/hr

Prabir Saha
10 years ago

36 KMPH

VEERU YERRAYYA
10 years ago

speed = (distance) / (time)

distance (d) = 100m

time (t) = 10sec = 10

speed = 100/ 0.1 = 1000m/s =  36km/h